OpenStudy (anonymous):

A neutral atom of sodium (Na) has an atomic number of 11 and a mass number of 23. How many neutrons would be found in the nucleus of this atom of sodium? 34 23 12 11

OpenStudy (anonymous):

atomic number is the number of protons and electrons and mass number is with neutrons... if you discard the mass of electron you can calculate number of neutrons by taking mass number and substracting number of protons (atomic number) so your answer is: 23-11 = ????

OpenStudy (anonymous):

Mass = Proton + neutrons

OpenStudy (anonymous):

Neutons = mass-Protons = 23 -11 = 12